A man weighing 60 kg is standing on a trolley weighing 240 kg.The trolley is resting on frictionless horizontal rails.If the man starts walking on the trolley along the rails at speed 1m/s,then after 4 seconds,his displacement relative to the ground will be

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
the trolley shall move backwards to conserve momentum. the back ward momentum would be shared by both the trolley and the man.
applying conservation of momentum
60*1 =  (240+60)*v
or
60 = 300 v
v = 1/5 m /sec
v = 0.2 m/sec
speed of man w.r.t ground = (1 – 0.2) m/sec
 = 0.8 m/sec
displacement of man = 0.8 * 4 = 3.2 m
 
Regards
Arun (askIITians forum expert)